OpenStudy (bloomlocke367):

How do I find if a matrix has an inverse? I know how to tell if two matrices are inverses, but not how to find if one has an inverse.

OpenStudy (rational):

you finished with probability !

OpenStudy (rational):

if a matrix has an inverse, then the determinant is not 0 if the determinant is not 0, then the matrix has an inverse

OpenStudy (bloomlocke367):

Yep! and matrices are kind of a review for me, but I don't have a calculator, and I don't know how to do this without a calculator.

OpenStudy (rational):

\[\text{determinant is not zero} \iff \text{invertible}\]

OpenStudy (rational):

whats teh exact complete question

OpenStudy (bloomlocke367):

does the matrix, A have an inverse? If so, what is \(A^{-1}\)?\[\left[\begin{matrix}-7 & -25 \\ 2 & 7\end{matrix}\right]\]

OpenStudy (rational):

do you know how to find the determinant ?

OpenStudy (bloomlocke367):

yes. ad-cb. where, \[\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\]

OpenStudy (rational):

Yes start by finding the determinant

OpenStudy (bloomlocke367):

1

OpenStudy (rational):

\(1 \ne 0\) so the inverse exists, go ahead and find it

OpenStudy (bloomlocke367):

how?

OpenStudy (rational):

http://gyazo.com/e69d2d115f0fc32ec73b4d509a6f88ba

OpenStudy (bloomlocke367):

so it's just \[\left[\begin{matrix}7 & 25 \\ -2 & -7\end{matrix}\right]\]

OpenStudy (bloomlocke367):

right?

OpenStudy (rational):

Yep! since the determinant is 1, dividing by 1 wont change the matrix